OpenStudy (anonymous):

I just need quick help. When you use the formula E=mc(theta), the change in temperature is given by (final temperature-initial temperature). Right?

OpenStudy (anonymous):

But I'm doing this question where ice is added to water at 28 degree Celsius and I'm required to calculate the final temperature of water. However, if I use (final temperature-initial temperature) I get the wrong answer.

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

well well ... yes

OpenStudy (experimentx):

but you forgot the latent heat of fusion.

OpenStudy (anonymous):

I did put the latent heat of fusion as well.

OpenStudy (experimentx):

or what do you call that ... melting!! also note the temperature of ice ... it has different heat sp. capacity.

OpenStudy (anonymous):

I used this equation: Heat lost by water at 28 degree Celsius= Heat gained by ice at 0 degree Celsius + Heat absorbed by water at 0 degree Celsius

OpenStudy (experimentx):

did change into cgs? or SI?

OpenStudy (anonymous):

\[m1c(T-28)=l _{f} m2 + m2c(T-0)\]

OpenStudy (anonymous):

Now if I use (28-T) in that equation, I get the right answer. Why is that so? Isn't the delta theta supposed to mean (final temp-initial temp)?

OpenStudy (experimentx):

of course 28-T is the delta(theta)

OpenStudy (experimentx):

the final temperature is T ... for initial temp for ice = 0 initial temp for water is 28

OpenStudy (anonymous):

That's what I'm saying. If T is the final temp for water, then shouldn't it be (T-28)? (Final-initial)

OpenStudy (experimentx):

delta theta for ice = T-0 delta theta for water = 28 -T

OpenStudy (experimentx):

yes yes ... you can do that, but you have to assume that heat of the system remains constant.

OpenStudy (experimentx):

do you get what I mean?

OpenStudy (anonymous):

Yes, the heat remains constant. But why (28-T)? Why not (T-28) as the formula states that delta theta is the final-initial temp?

OpenStudy (experimentx):

if you put T-28, you have to assume that heat remains constant. change in heat of water + change in heat of ice + latent heat of fusion = constant.

OpenStudy (experimentx):

the constant = 0 I mean the change in heat of the whole system is constant. just consider the direction of heat flow.

OpenStudy (experimentx):

if heat is lost then 28-T if heat is gained, then T-28

OpenStudy (anonymous):

The simple answer is 28 degrees minus 0 degrees equals 28 degrees as the change in temperature. The direction of change is 28 degrees cooler. As long as there is ice to melt the temperature will drop to 0 degrees. You only need to use the heat of fusion to determine how much ice needs to melt to bring the water down to 0 degrees, a different problem.